OpenStudy (anonymous):

An auditorium earned $25,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

if the number of back seat is say \(x\) then the number of front seats is \(2x\) therefore \(50x+75\times 2x=25000\) or \[50x+150x=25000\] or \[200x=25000\]

OpenStudy (anonymous):

solve for \(x\) but then make sure to double the answer because it asks for "front seats"

OpenStudy (anonymous):

125 then double the answer to 250?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

thank you so much!